The total market value of Okefenokee Real Estate Company’s equity is $6 million, and the total value of its debt is $4 million. The treasurer estimates that the beta of the stock currently is 1.5 and that the expected risk premium on the market is 10%. The Treasury bill rate is 6%, and investors believe that Okefenokee's debt is essentialy free of default risk.

a. What is the required rate of return on Okefenokee stock? (Do not round intermediate calculations. Enter your answer as a whole percent.)
  Required rate of return %  
b.

Estimate the WACC assuming a tax rate of 40%. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

  WACC %  
c.

Estimate the discount rate for an expansion of the company’s present business. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

  Discount rate %  
d.

Suppose the company wants to diversify into the manufacture of rose-colored glasses. The beta of optical manufacturers with no debt outstanding is 1.7. What is the required rate of return on Okefenokee’s new venture? (You should assume that the risky project will not enable the firm to issue any additional debt.) (Do not round intermediate calculations. Enter your answer as a whole percent.)

  Required rate of return %  
Answer #1
a) Required rate of return on stock per CAPM = 6%+1.5*10% = 21.00%
b) WACC = 6%*(1-40%)*4/10+21%*6/10 = 14.04%
c) Discount rate of expansion in present business = WACC = 14.04%
d) As no more debt is possible, the cost of unlevered equity of the industry is the required return for the new venture = 6%+1.7*10% = 23.00%